OpenStudy (anonymous):

5^n+1 is divisible by 6

ganeshie8 (ganeshie8):

are you trying to prove that statement ? please provide full details

OpenStudy (lieutenantgeneral):

false + 5^n

OpenStudy (lieutenantgeneral):

www.wolframalpha.com

OpenStudy (lieutenantgeneral):

http://www.wolframalpha.com/input/?i=5^n%2B1+is+divisible+by+6

OpenStudy (anonymous):

no divisible. 5^2+1=25+1=26 is not divisible by 6

OpenStudy (anonymous):

*not

OpenStudy (anonymous):

The full thing is: Which of the following is a counterexample for the given statement. 5^n+1 is divisible by 6. a. n=1 c. n=3 b. n=2 d. n=5

OpenStudy (anonymous):

b. n=2

OpenStudy (anonymous):

How do you know that?

OpenStudy (anonymous):

5^2+1=25+1=26 is not divisible by 6 @xxcamxx

OpenStudy (anonymous):

Ohhhh okay! Thank you